# Top-Theta Attention: Sparsifying Transformers by Compensated Thresholding

### v1

@@ Abstract
The attention mechanism is essential for the impressive capabilities of transformer-based Large Language Models (LLMs). However, calculating attention is computationally intensive due to its quadratic dependency on the sequence length. We introduce a novel approach called Top-Theta Attention, or simply Top-{greek}j, which selectively prunes less essential attention elements by comparing them against carefully calibrated thresholds. This method greatly improves the efficiency of self-attention matrix multiplication while preserving model accuracy, reducing the number of required V cache rows by 3x during generative decoding and the number of attention elements by 10x during the prefill phase. Our method does not require model retraining; instead, it requires only a brief calibration phase to be resilient to distribution shifts, thus not requiring the thresholds for different datasets to be recalibrated. Unlike top-k attention, Top-{greek}j eliminates full-vector dependency, making it suitable for tiling and scale-out and avoiding costly top-k search. A key innovation of our approach is the development of efficient numerical compensation techniques, which help preserve model accuracy even under aggressive pruning of attention scores.
Machine Learning, ICML
@@ 1 Introduction
The transformer architecture has revolutionized both natural language processing (Vaswani, 2017) and computer vision (Dosovitskiy, 2020) by enabling models to capture complex dependencies through self-attention mechanisms effectively. However, despite its advantages, the attention mechanism suffers from quadratic time and linear memory complexity (Keles et al., 2023). Furthermore, the commonly used key-value (KV) cache optimization increases the memory requirement linearly with sequence length during the generative decoding phase. As a result, cache size requirements often exceed the physical limits of available memory (Ge et al., 2024), and memory bandwidth becomes a major bottleneck.
Attention approximations (Wang et al., 2024; Fuad & Chen, 2023), such as sparsification, promise a solution to these challenges by reducing the number of computations and the amount of data moved, focusing only on the most relevant tokens in a sequence. Research on sparsification has been predominantly focused on either fixed-sparsity patterns, which assume that specific token locations in the sequence are always more important (e.g., the first and last tokens in the sequence), or the content-based sparsity patterns, which require evaluating the attention scores to decide which tokens are more important (e.g., Top-k attention (Gupta et al., 2021)).
Our work focuses on the more challenging, content-based sparsity patterns. To exploit the sparsity potential, we investigate the sparsification of the attention elements through pruning by comparing to a threshold value. We calibrate the thresholds to select a desired average number of $k$ important tokens in every attention row. We find that calibrating model-specific thresholds is sufficient to replace the top-k search over the attention elements. Once the important tokens have been quickly determined by thresholding, the remaining tokens can be excluded from participating in the softmax computation and in the multiplication by the V-matrix, thus avoiding the need to load them. Moreover, to preserve the high accuracy in the downstream task, we propose numerical compensation methods such as softmax denominator compensation and mean V-row compensation.
@@
The goal of this work is to unlock the potential of the content-based sparsity for more compute- and memory-efficient attention while preserving a high-quality output. Our fundamental finding is:
@@
Based on our fundamental finding, we derive the Top-{greek}j attention method, unlocking the following advantages:
@@
    Efficiency – 3× fewer V-rows needed and 10×  less attention elements needed for LLaMA2 and LLaMA3 models to achieve same accuracy.
@@
    No retraining is needed to calibrate the thresholds, which takes a few hundred calibration samples.
@@
    Distribution shift resilience – we observe that the same thresholds of a given model apply across different input domains. In other words, thresholds are a strong characteristic of the model; hence, the calibration is required only once per model.
Paper organization
Section 2 recaps the self-attention mechanism, emphasizing efficient inference based on sparsity and the Top-k method.  Section 3 describes our Top-{greek}j method and shows how it overcomes the limitations of Top-k and our numerical compensations that further improve the accuracy of sparsification. Section 4 first evaluates our method, including all compensations, and further provides ablations of its different aspects. Sections 5 and 6 review relevant studies and summarize our findings.
@@ 2.1 Transformer Models
Transformer-based models consist of layers that process token sequences. Most current large language models (LLMs) for text generation are autoregressive decoder-only transformer layers, known for their strong zero-shot generalization (Wang et al., 2022) and use in chatbots and productivity tools. These models first process the entire input sequence in a single forward pass (prefill phase) and then generate additional tokens one at a time, each requiring a separate forward computation (generative decoding phase). This research aims to enhance the decoder-only transformers.
@@ 2.2 Self-Attention and Sparsity
Multi-head self-attention (MHA) is the first computational step of the transformer layer. The MHA receives as an input a sequence of tokens $X\in\mathbb{R}^{n\times D}$ where $n$ is the sequence length and $D$ is the hidden dimension. Each of the heads processes $X$ in parallel by first multiplying it by 3 different trained matrices $W_{Q},W_{K},W_{V}\in\mathbb{R}^{D\times d}$ obtaining 3 matrices $Q,K,V\in\mathbb{R}^{n\times d}$ and adding a positional encoding (e.g., RoPE (Su et al., 2024)) to them.
$Q=XW_{Q},K=XW_{K},V=XW_{V}$ (1)
Then, a pre-softmax attention matrix $A$ is computed from matrices $Q$ and $K$ (2). Although $A$ matrix is often normalized by $\sqrt{d}$ for numerical stability and masked by a causality mask, we omit these 2 steps for simplicity.
$A=QK^{T}$ (2)
After that, each row of the $A$ matrix is normalized by the Softmax operation yielding the post-softmax attention matrix $S$ (3) followed by a multiplication by $V$ (4).
$S$ $=\mathrm{row\_softmax}(A)$ (3)
$P$ $=SV$ (4)
@@
Both pre-, and post-softmax matrices are of the shape $n\times n$, in which the element $(i,j)$ element signifies the importance of token $j$ for token $i$. The pre-softmax attention $A$ has a range of all real numbers distributed normally, whereas the post-softmax $S$ has the range of $[0,1]$ and is distributed log-normally. The attention elements in the initial layers exhibit a less skewed distribution (i.e., high entropy), whereas the following layers have a more concentrated distribution (i.e., low entropy) with a few rare high-attention values (Vig & Belinkov, 2019; Nahshan et al., 2024). In both $A$ and $S$, a lower attention value has a lower contribution to the further computation because $S$ is obtained through an order-preserving transformation of $A$, after which $S$ is multiplied by the matrix $V\in\mathbb{R}^{n\times d}$ resulting in the output matrix $P\in\mathbb{R}^{n\times d}$ (4). Therefore, small values in column $i$ in $S$ diminish the impact of the row $i$ of the $V$ matrix. This fundamental property of the attention elements allows ranking them according to their significance and pruning the least significant ones for sparsification.
@@
MHA employs a well-established performance optimization called KV-cache (Shi et al., 2024) which allows processing a single embedded token $\bm{x}\in\mathbb{R}^{D}$, computing only the current token’s $\bm{q},\bm{k},\bm{v}\in\mathbb{R}^{d}$ vectors, while the complete $K,V$ matrices are loaded from the cache (avoiding recomputation) and the new $\bm{k},\bm{v}$ vectors are appended to them. In this situation, the attention matrices simplify to vectors $\bm{a},\bm{s}\in\mathbb{R}^{n}$ representing the attention of the currently generated token to all the previous tokens, and the $P$ matrix becomes a single token embedding $\bm{p}=\bm{s}V\in\mathbb{R}^{d}$. Due to the large size of the KV-caches, especially as the sequence length grows longer, the computation of the self-attention during the generative decoding is heavily limited by memory bandwidth, dominated by loading all the $n$ rows of the K and V matrices, while only performing 1 multiply-add per value read from memory. Due to the memory bottleneck, a reduction of memory reads directly leads to a corresponding speed up.
An established variant of multi-head self-attention is called grouped query attention (GQA). The GQA introduces sharing a single pair of $K,V$ matrices for a group of $g$ heads (queries), which reduces the amount of $K,V$ data to load by a factor of $g$. Recent studies have shown that GQA has only a marginal impact on the accuracy of the downstream task, making it a favorable optimization (Ainslie et al., 2023).
@@ 2.3 Top-k Attention
One widely adopted approach of sparsifying the attention row is finding its top $k$ out of $n$ elements and discarding the rest, where $k$ is a hyperparameter (Gupta et al., 2021). However, since the attention rows are often partitioned (tiling across the sequence dimension) and processed in parallel, computing the exact top-k values in a vector imposes an undesired full-row dependency, thereby constraining the partitioning strategies. Moreover, the computation of the top elements requires several computational steps (Zhang et al., 2023), leading to a logarithmic complexity at the very best case. We conjecture that the top-k search algorithms can be replaced by comparison to a carefully calibrated threshold.
@@ 3 Top-{greek}j Method
Figure 1: Top-{greek}j attention inference at generative decoding.
Top-{greek}j attention involves comparing each attention vector against a calibrated threshold. Attention elements that fall below the threshold are pruned away from subsequent computations, enhancing the efficiency and focus of the model. Our underlying assumption is that a particular distribution of values characterizes each row of the attention matrix, therefore using a static threshold should keep approximately the desired number of selected attention elements. The benefit of using a threshold instead of computing the top-k attention values is that thresholding is a simple elementwise operation, requiring only a constant time and involving no row dependency. In contrast, ranking the elements of a vector requires at least logarithmic time and depends on the full length of the vector. Top-{greek}j can be seen as an approximation of Top-k attention, as it allows calibrating the thresholds for specific user-defined $k$.
The rest of this section provides a detailed discussion of (1) the calibration process for the thresholds and (2) their application for efficient inference.
@@ 3.1 Threshold Calibration
Given a transformer model, a set $\mathcal{C}$ of calibration data samples, and the desirable number $k$ of elements to keep per attention row, we propose to calibrate a single scalar threshold $\theta_{l,h,r}(k)\in\mathbb{R}$ for every transformer layer $l$, head $h$, and attention row id $r$. The objective of the threshold calibration is to determine a threshold value $\theta_{l,h,r}(k)$ such that on average $k$ elements are greater than $\theta_{l,h,r}(k)$. The calibration procedure listed in Algorithm 1 calibrates the thresholds of a single attention head in a single layer, thus it can be launched as a part of the forward computation of self-attention. The key step in the calibration is collecting the $(\frac{r-k}{k})$-quantile of the $r^{\mathrm{th}}$ attention row values for each calibration sample into a set $\Theta_{r}$ (i.e., l. 7). Each quantile is essentially a per-calibration sample threshold. After collecting the per-sample thresholds, the final threshold for attention row $r$ is computed as an average over $\Theta_{r}$ (l. 19). In  Section 4.4, we show the importance of calibrating individual thresholds for every row id. To obtain the thresholds for the entire model, Algorithm 1 should be applied independently to all other layers and heads in the model.
Algorithm 1 Calibrate($\mathcal{C},k,\alpha$) - 1 head threshold calibration
0:  $\mathcal{C}$ (Calibration set of inputs)
0:  $k\in\mathbb{N}$ (elements to keep per attention row)
0:  $\alpha\in\mathbb{R}$ (calibration offset in standard deviations)
1:  $\Theta_{r}=\emptyset,\forall r$empty sets of observed thresholds}
2:  for $X\in\mathcal{C}$ do
3:     if $\mathrm{is\_prefill}(X)$ then
4:        $A=Q(X)K^{T}(X)$
5:        $n=\mathrm{num\_rows}(A)$
6:        for $r=k$ to $n-1$ do
7:           $\Theta_{r}$ = $\Theta_{r}\cup\{\mathrm{quantile}_{\frac{n-k}{n}}(A_{r})\}$
8:           $A_{r}=\mathrm{top}_{k}(A_{r})$
@@
Calibration hyperparameters
The main parameter is $k$, controlling the target number of elements to keep per attention row. Generally, $k$ can be individually chosen for every layer, head, and token position. To reduce the hyperparameterization, we found that selecting a single $k$ for all heads and attention row indices within the given layer is sufficient to preserve the quality of results. However, it is necessary to allow the flexibility of choosing different $k$ values for the different layers due to their inherently different sparsity levels. The hyperparameter $\alpha$ is used to offset the threshold to a higher value when keeping the target $k$ elements should be followed more conservatively (l. 19 in Alg. 1). See Figure 2 for the visualization of the threshold values collected for a specific layer, head, row, and selected threshold.
@@
Figure 2: Distribution of the set $\Theta_{1000}$, collected during calibration of threshold $\theta_{10,20,800},\alpha=0.1$ on Hellaswag, LLaMA2-7b model.
Calibration details

### v2

@@ Abstract
We present Top-Theta (Top-{greek}j) Attention, a training-free method for sparsifying transformer attention during inference. Our key insight is that static, per-head thresholds can be calibrated to retain the desired constant number of significant elements per attention row. This approach enables content-based sparsity without retraining, and it remains robust across data domains. We further introduce compensation techniques to preserve accuracy under aggressive sparsification, establishing attention thresholding as a practical and principled alternative to top-k attention. We provide extensive evaluation on natural language processing tasks, showing that Top-{greek}j achieves 3–10× reduction in $\bm{V}$-cache usage and up to 10× fewer attention elements during inference while degrading no more than 1% in accuracy.
@@ 1 Introduction
The transformer architecture has revolutionized natural language processing Vaswani et al. (2017) and computer vision Dosovitskiy (2020) by enabling models to capture complex dependencies through self-attention mechanisms effectively. However, despite its advantages, the attention mechanism suffers from quadratic time and linear memory complexity Keles et al. (2023). Furthermore, the commonly used key-value (KV) cache optimization increases the memory requirement linearly with sequence length during the generative decoding phase. As a result, cache size requirements often exceed the physical limits of available memory Ge et al. (2024), and memory bandwidth becomes a major bottleneck.
Attention approximations Wang et al. (2024); Fuad and Chen (2023), such as sparsification, promise a solution to these challenges by reducing the number of computations and the amount of data moved, focusing only on the most relevant tokens in a sequence. Research on sparsification has been predominantly focused on either fixed-sparsity patterns, which assume that specific token locations in the sequence are always more important (e.g., the first and last tokens in the sequence), or the content-based sparsity patterns, which require evaluating the attention scores to decide which tokens are more important (e.g., Top-k attention by Gupta et al. (2021)).
Our work focuses on the more challenging, content-based sparsity patterns. To exploit the sparsity potential, we investigate the sparsification of the attention elements through pruning by comparing to a threshold value. We calibrate the thresholds to select a desired average number of $k$ important tokens in every attention row. We find that calibrating model-specific thresholds is sufficient to replace the top-k search over the attention elements. Once the important tokens have been quickly determined by thresholding, the remaining tokens can be excluded from participating in the softmax computation and in the multiplication by the $\bm{V}$-matrix, thus avoiding the need to load the corresponding $\bm{V}$-rows. Moreover, to preserve the high accuracy in the downstream task, we propose numerical compensation methods such as softmax denominator compensation and mean $\bm{V}$-row compensation.
@@
This work is taming the potential of content-based sparsity for more compute- and memory-efficient attention with negligible accuracy degradation. Our fundamental finding is:
@@
From this fundamental finding, we derive the Top-{greek}j attention method, unlocking the following advantages:
@@
    Efficiency. 3× to 10× fewer $\bm{V}$-rows needed and 10×  less attention elements needed for LLaMA2 and LLaMA3 models to achieve same accuracy.
@@
    No retraining. Threshold calibration requires only a few hundred samples and no retraining.
@@
    Distribution shift resilience. Thresholds remain consistent across input domains, representing a fundamental model characteristic that requires only one-time calibration.
@@ 2.1 Transformer Models
Modern Large Language Models (LLM) used for text generation primarily employ decoder-only transformer layers, noted for strong zero-shot generalization Wang et al. (2022) and widespread success in chatbots and productivity tools. These models operate in two phases: processing the entire input at once (prefill) and generating tokens sequentially (generative decoding). Our research aims to enhance the underlying self-attention mechanism of decoder-only transformers.
@@ 2.2 Self-Attention and Sparsity
Multi-head self-attention (MHA) is the first computational step of the transformer layer. The MHA receives as input a sequence of tokens $\bm{X}\in\mathbb{R}^{n\times D}$ where $n$ is the sequence length and $D$ is the hidden dimension. Each of the heads processes $\bm{X}$ in parallel by first multiplying it by 3 different trained matrices $\bm{W_{Q}},\bm{W_{K}},\bm{W_{V}}\in\mathbb{R}^{D\times d}$ obtaining 3 matrices $\bm{Q},\bm{K},\bm{V}\in\mathbb{R}^{n\times d}$ and adding a positional encoding (e.g., RoPE Su et al. (2024)) to them. Then, a pre-softmax attention matrix $\bm{A}$ is computed from matrices $\bm{Q}$ and $\bm{K}$ (1). Although $\bm{A}$ matrix is often normalized by $\sqrt{d}$ for numerical stability and masked by a causality mask, we omit these 2 steps for simplicity.
$\bm{A}=\bm{QK}^{T}$ (1)
After that, each row of the $\bm{A}$ matrix is normalized by the Softmax operation, yielding the post-softmax attention matrix $\bm{S}$ (2), which is then multiplied by $\bm{V}$ (3).
$\bm{S}$ $=\mathrm{row\_softmax}(\bm{A})\triangleq\left[\frac{e^{A_{ij}}}{\sum_{k=1}^{n}e^{A_{ik}}}\right]_{\begin{subarray}{c}1\leq i\leq n\\ 1\leq j\leq n\end{subarray}}$ (2)
$\bm{P}$ $=\bm{S}\bm{V}$ (3)
@@
Both pre-, and post-softmax matrices are of the shape $n\times n$, in which the element $(i,j)$ signifies the importance of token $j$ for token $i$. The pre-softmax attention $\bm{A}$ has a range of all real numbers distributed normally, whereas the post-softmax $\bm{S}$ has the range of $[0,1]$ and its distribution resembles log-normal, with the majority of the values concentrated near 0. The attention elements in the initial layers exhibit a less skewed distribution (i.e., high entropy), whereas the following layers have a more concentrated distribution (i.e., low entropy) with a few rare high-attention values Vig and Belinkov (2019); Nahshan et al. (2024). In both $\bm{A}$ and $\bm{S}$, a lower attention value has a lower contribution to the further computation because $\bm{S}$ is obtained through an order-preserving transformation of $\bm{A}$, after which $\bm{S}$ is multiplied by the matrix $\bm{V}\in\mathbb{R}^{n\times d}$ resulting in the output matrix $\bm{P}\in\mathbb{R}^{n\times d}$ (3). Therefore, small values in column $i$ in $\bm{S}$ diminish the impact of row $i$ of the $\bm{V}$ matrix. This fundamental property of the attention elements allows ranking them according to their significance and pruning the least significant ones for sparsification.
@@
MHA employs a well-established performance optimization called KV-cache Shi et al. (2024) which allows processing a single embedded token $\bm{x}\in\mathbb{R}^{D}$, computing only the current token’s $\bm{q},\bm{k},\bm{v}\in\mathbb{R}^{d}$ vectors, while the complete $\bm{K},\bm{V}$ matrices are loaded from the cache (avoiding recomputation) and the new $\bm{k},\bm{v}$ vectors are appended to them. In this situation, the attention matrices simplify to vectors $\bm{a},\bm{s}\in\mathbb{R}^{n}$ representing the attention of the currently generated token to all the previous tokens, and the $\bm{P}$ matrix becomes a single token embedding $\bm{p}=\bm{s}\bm{V}\in\mathbb{R}^{d}$. Due to the large size of the KV-caches, especially as the sequence length grows longer, the computation of the self-attention during the generative decoding is heavily limited by memory bandwidth, dominated by loading all the $n$ rows of the $\bm{K}$ and $\bm{V}$ matrices, while only performing 1 multiply-add per value read from memory. Due to the memory bottleneck, a reduction of memory reads directly leads to a corresponding speed up.
A variant of multi-head self-attention is called grouped query attention (GQA). The GQA introduces sharing a single pair of $\bm{K},\bm{V}$ matrices for a group of $g$ heads (queries), which reduces the amount of $\bm{K},\bm{V}$ data to load by a factor of $g$. Studies have shown that GQA has a marginal impact on the downstream task accuracy, making it a favorable optimization Ainslie et al. (2023).
@@ 2.3 Top-k Attention
One widely adopted approach to sparsifying the attention row is finding its top $k$ out of $n$ elements and discarding the rest, where $k$ is a hyperparameter Gupta et al. (2021). However, since the attention rows are often partitioned (tiling across the sequence dimension) and processed in parallel, computing the exact top-k values in a vector imposes an undesired full-row dependency, thereby constraining the tiling strategies. Moreover, the computation of the top elements requires several computational steps Zhang et al. (2023), leading to a logarithmic complexity at the very best case. We conjecture that the top-k search algorithms can be replaced by comparison to a carefully calibrated threshold.
@@ 3 Top-{greek}j Method
Figure 1: Two variants of Top-{greek}j attention for inference at generative decoding.
Our proposed Top-{greek}j attention involves comparing each attention vector against a calibrated threshold. Attention elements that fall below the threshold are pruned away from subsequent computations, enhancing the efficiency and the focus of the model. Our underlying assumption is that a particular distribution of values characterizes each row of the attention matrix; therefore, using a static threshold should keep approximately the desired number of selected attention elements. Our motivation for using a threshold-based method instead of computing the top-k attention values is that thresholding is a simple elementwise operation, requiring only a constant time and involving no row dependency. In contrast, ranking the elements of a vector requires at least logarithmic time and depends on the full length of the vector. Top-{greek}j can be seen as an approximate Top-k, as it allows calibrating the thresholds for specific user-defined $k$.
@@ 3.1 Threshold Calibration
Algorithm 1 Calibrate($\mathcal{C},k,\alpha$) - 1 head threshold
[…]
2:for $\bm{X}\in\mathcal{C}$ do
3:  if $\mathrm{is\_prefill}(\bm{X})$ then
4:   $\bm{A}=\bm{Q}(\bm{X})\bm{K}^{T}(\bm{X})$
5:   $n=\mathrm{num\_rows}(\bm{A})$
6:   for $r=k$ to n1n-1 do
7:    $\Theta_{r}$ = $\Theta_{r}\cup\{\mathrm{quantile}_{\frac{n-k}{n}}(\bm{A}_{r})\}$
8:    $\bm{A}_{r}=\mathrm{top}_{k}(\bm{A}_{r})$
@@
Prior to using Top-{greek}j for efficient inference, its thresholds need to be calibrated with respect to a user-defined parameter $k$. To this end, we present Algorithm˜1 that calibrates a threshold $\theta_{r}(k)\in\mathbb{R}$ for attention row id $r$ of a given transformer layer and head. The calibration is performed offline, before the model deployment, by collecting the $(\frac{r-k}{k})$-quantile of each attention row from a set of calibration samples (l. 7), then averaging these per-sample thresholds to obtain $\theta_{r}(k)$ (l.19). This algorithm has to be performed for all layers and heads in parallel, and can be integrated into a single forward pass of the model (then disabled after the calibration is complete). We found that using a single $k$ per layer suffices for robust performance, though different $k$ can be chosen per head or row.
For practical calibration, we recommend a calibration set of a few hundred samples, as we found that increasing the calibration set size mainly improves the fidelity of thresholding to keep the desired $k$ elements per row, but does not benefit the accuracy of the downstream task (see Appendix˜B). An optional offset hyperparameter $\alpha$ allows conservative adjustment of the final threshold. See Figure˜2 for the visualization of the threshold values collected for a specific layer, head, row, and selected threshold.
@@
Figure 2: Distribution of the set $\Theta_{1000}$, during calibration of threshold $\theta_{10,20,800},\alpha=0.1$ on Hellaswag, LLaMA2-7b.
During calibration, a “Top-k at calibration” step (l. 8,15) ensures that subsequent layers’ activations reflect the sparsification pattern, improving threshold stability at test time. We note that Algorithm˜1 in its presented form is calibrating pre-softmax thresholds, whereas it can be adapted to post-softmax by taking the quantile on the post-softmax attention $\bm{S},\bm{s}$ rather than on the $\bm{A},\bm{a}$. We also note that calibrating for rows $r<k$ is unnecessary for the first $k$ rows due to causal masking.
Once the calibration is complete, the calibrated thresholds can be stored along with the model parameters, as they take negligible memory. For example, when 1200 per-row float16 thresholds are calibrated (as in ARC-C) for every head, the total memory for thresholds for an entire LLaMA-3-70B model (80 layers, 64 heads) is only $80\cdot 64\cdot 1200\cdot 2=11.8$ Mbytes (0.0008% of the model size).
Figure˜3 visualizes two threshold variants of the 11th layer of LLaMA2-7b, which were calibrated for $k=64$: pre-softmax, and post-softmax. Notably, the threshold values depend not only on the sequence length (x-axis) but also on the attention head (different colors), justifying the calibration of an individual threshold for each attention head. Secondly, thresholds obtained pre-softmax approach a constant value as the sequence length increases, whereas thresholds obtained from the post-softmax tend to decrease with longer sequence lengths. The latter is an effect of the softmax normalization that reduces the share of essential tokens from the total row sum of 1 as the sequence length increases; therefore, keeping them requires lowering the threshold. The scattering artifact of the thresholds in the longer calibrated sequence lengths is due to the scarcity of the calibration samples in the dataset that corresponded to these sequence lengths (low purple bars in Figure˜3).  Appendix˜C presents a visualization of thresholds from additional layers.
Threshold function fitting
Our main proposal is to deploy the calibrated thresholds as complementary model parameters (one threshold parameter per every layer, head, and sequence length). However, as can be observed in Figure˜3, the thresholds have a rather regular behavior as a function of the sequence length (i.e., of the attention row id). Therefore, we can fit a function to it and thereby parameterize hundreds of thresholds by a few parameters. Such a fitting can be obtained via a weighted least squares solution, with the weights being the number of calibration samples obtained for each sequence length. With this weighing, the scattering artifact observed in the longer sequence lengths will be mitigated.
Multi-k calibration
It is possible to calibrate multiple threshold sets $\theta_{l}(k_{1}),\theta_{l}(k_{2}),$... targeted at different number of top elements $k_{1},k_{2},$... to be preserved. Following such a calibration, one can dynamically choose the desired accuracy-performance tradeoff at inference by applying a set of thresholds with a different $k$, a desired capability for a flexible LLM serving Miao et al. (2023). Instead of applying  Algorithm˜1 several times with the different $k$ values, we have developed a generalized calibration algorithm, which performs a single pass over the calibration samples. The generalized procedure collects up to $n$ thresholds for every attention vector of length $n$ in the calibration sample, merges this threshold information, and eventually outputs for each (layer, head, row) a function $\theta(k)$ from $k$ to a threshold that satisfies it. Although such a calibration procedure poses large memory requirements, reducing its intensity by sampling fewer than $n$ thresholds per vector is possible (see Appendix˜D).
3.2 Top-{greek}j Attention Inference
The Top-{greek}j attention inference operation in Figure˜1 focuses on a single transformer layer $l$, single head $h$, and single attention row $n$ of length $n$ (as in generative decoding). The attention vector is compared against the calibrated threshold value $\theta_{n}$. Attention elements that do not pass the threshold are discarded. Instead of multiplying by the entire $\bm{V}$ matrix, only the selected $\tilde{k}$ row indices are loaded to a compact matrix $\tilde{\bm{V}}$, which is used to compute the final product $\bm{p}$.
Technical details: (i) Since the calibration set might not have covered the entire range of sequence lengths $r\in\{k,k+1,\ldots\}$, at inference, the threshold of the nearest calibrated sequence length is used. (ii) Threshold calibrated for a target $k$ is guaranteed to select $k$ attention elements per row only on average, i.e., actual counts ($\tilde{k}$) may slightly vary due to input-dependent attention distributions. To handle cases where $\tilde{k}>k$, we experimented with capping the selection at $k$ elements, prioritizing first and last tokens, but found that this noticeably degraded performance on generative tasks. Instead, we mitigated the variability of $\tilde{k}$ by increasing the calibration set size (Appendix˜B).
The rest of this section introduces two efficient compensation mechanisms that mitigate accuracy degradation in Top-{greek}j attention through improved mathematical approximation of full self-attention.
3.2.1 Softmax Denominator Compensation (SDC)